Circular No. 8504

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 G1 (LINEAR) An object reported as apparently asteroidal by the LINEAR search

program (discovery observation given below), and posted on the ‘NEO Confirmation Page’, has been found to show a 9′′ coma and an 11′′ tail in p.a. 180^{o} on R-band images taken by C. Hergenrother with the 1.54-m Kuiper reflector at Catalina on Apr. 4.45 UT.

2005 UT  α_{2000}  δ_{2000}  Mag.
Apr. 1.39390  18^{h}12^{m}30.^{s}94  +46^{o}18′37′′.3  19.1

The available astrometry (including prediscovery LINEAR observations from Mar. 22), the preliminary parabolic orbital elements (T = 2006 Feb. 28.5 TT, q = 4.95 AU, ω = 114^{o}, Ω = 300^{o}, i = 108^{o}, equinox 2000.0), and an ephemeris appear on MPEC 2005-G23.

SUPERNOVAE 2005az, 2005bb, AND 2005bc Further to IAUC 8493, J. Burket, H. Pugh, and W. Li report the

LOSS discovery of two apparent supernovae on unfiltered KAIT images. Further to IAUC 8499, T. Puckett and L. Cox also report their independent discovery of SN 2005bc. Tabulated LOSS discovery data:

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005bb Apr. 1.40 12^{h}57^{m}11.^{s}87 − 1^{o}42′17′′.4 19.8 4′′.1 W, 6′′.9 N
2005bc Apr. 2.44 14 37 15.05 +38 27 23.1 16.6 4′′.6 E, 7′′.5 N

Additional magnitudes, from KAIT unless otherwise noted: SN 2005bb in UGC 8067, 2004 Dec. 25.41 UT, [20.5; 2005 Jan. 21.44, [19.0; Feb. 3.41, [17.5; Mar. 6.40, 18.0; Apr. 2.41, 19.0. SN 2005bc in NGC 5698, Mar. 11.43, [20.0; Apr. 4.39, 16.4 (Puckett and Cox). A Sloan Digital Sky Survey image of UGC 8067 from 2000 Feb. 12.46 showed nothing at the position of SN 2005bb (limiting mag 22.2). Puckett reports position end figures 15^{s}.05, 23′′.7 for SN 2005bc.

R. Quimby, P. Hoeflich, and J. C. Wheeler, University of Texas; and C. Gerardy, Imperial College, report that a spectrogram (range 430–890 nm) of SN 2005az (cf. IAUC 8503), obtained on Apr. 2.42 UT with the 9.2-m Hobby/Eberly Telescope (+ Marcario Low-Resolution Spectrograph) by M. Shetrone and E. Terrazas, shows it to be a type-Ic supernova; the spectrum resembles that of SN 1994I near maximum light (e.g., Clocchiatti et al. 1996, Ap.J. 462, 462).

2005 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green